The intensity distribution of solar radiation peaks at a wavelength of approximately 500 nm. Assuming the Sun radiates like a black body, calculate the flux of photons leaving the sun in W/m^2

.
1
Expert's answer
2020-08-30T02:22:59-0400

The flux of photons (j) is found from Stefan Boltzmann law: 

j = σT4, where constant σ = 5.67*10-8 W/(m2*K4)

T is found from Wien displacement law: 

λmax = σw /T, where constant σw = 0.002898 m*K

λmax = 5*10-7 m = 0.002898/T, hence, T = 5796 K

The flux of photons is:

j = 5.67*10-8*(5796)4 = 6.4*10W/m2